# Explicit Second-Order Min-Max Optimization Methods with Optimal Convergence Guarantee

### v4

@@ 5 Numerical Experiments
We evaluate the performance of our methods for min-max optimization problems with synthetic and real data. The baseline methods include extragradient and optimistic gradient descent ascent method (we refer to them as EG1 and OGDA1), stochastic variants of EG and OGDA (which we refer to as SEG and SOGDA) and second-order variants of EG and OGDA (which we refer to as EG2 and OGDA2). We exclude the methods in Huang et al. (2022) since their guarantees are proved under additional error bound conditions. All methods were implemented using MATLAB R2023b on a MacBook Pro with an Intel Core i9 2.4GHz and 16GB memory.
@@ 5.1 Cubic regularized bilinear min-max problem
where $\rho>0$, the entries of $\mathbf{b}\in\mathbb{R}^{n}$ are generated independently from $[-1,1]$ and $A\in\mathbb{R}^{n\times n}$ is given by
$A=\begin{bmatrix}1&-1&&&\\ &1&-1&&\\ &&\ddots&\ddots&\\ &&&1&-1\\ &&&&1\end{bmatrix}.$
This min-max optimization problem is convex-concave and the function $f$ is $\rho$-Hessian Lipschitz. It has a unique global saddle point $\mathbf{x}^{\star}=A^{-1}\mathbf{b}$ and $\mathbf{y}^{\star}=-\frac{\rho}{2}\|\mathbf{x}^{\star}\|(A^{\top})^{-1}\mathbf{x}^{\star}$. We use the restricted gap function defined in Section 2 as the evaluation metric. In our experiment, the parameters are chosen as $\rho=\frac{1}{20n}$ and $n\in\{50,100,200\}$. Since the exact second-order information is available here, we use Algorithm 2 with $\tau_{0}=0$ (the exact Hessian) and compute each subproblem up to a very high accuracy with $\kappa_{m}=10^{-8}$ using our subroutine. The baseline methods include EG1, OGDA1, EG2 and OGDA2 where all of these methods require the exact second-order information. We implement EG2 using the pseudocode of Bullins and Lai (2022, Algorithm 5.2 and 5.3) with the fine-tuning parameters. The implementation of OGDA2 is based on the code provided by the author of Jiang and Mokhtari (2022) with the line search parameter $(\alpha,\beta)=(0.5,0.8)$.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 1: Performance of all the algorithms for $n\in\{50,100,200\}$ when $\rho=\frac{1}{20n}$ is set. The numerical results are presented in terms of iteration count (Top) and computational time (Bottom).
Figure 1 highlights that the second-order methods can be far superior to the first-order methods in terms of solution accuracy: the first-order methods barely make any progress when the second-order methods converge successfully. Moreover, our method outperforms other second-order methods (Bullins and Lai, 2022; Jiang and Mokhtari, 2022) in terms of both iteration numbers and computational time thanks to its simple scheme without line search. Nonetheless, we remark that this does not eliminate the advantages of using line search in min-max optimization. Indeed, we find that the line search scheme from Jiang and Mokhtari (2022) is very powerful in practice and their methods with aggressive choices of $(\alpha,\beta)$ outperforms our method in many cases. However, such choices make their method unstable. Thus, we choose $(\alpha,\beta)=(0.5,0.8)$ which is conservative yet more robust. We believe it is promising to study the line search scheme from Jiang and Mokhtari (2022) and see if modifications can speed up second-order min-max optimization in a universal manner.
Table 1: Statistics of datasets for AUC maximization.
Name Description $N$ $n$ Scaled Interval
a9a UCI adult 32561 123 $\left[0,1\right]$
covetype forest covetype 581012 54 $\left[0,1\right]$
w8a - 49749 300 $[0,1]$
@@ 5.2 AUC maximization problem
The problem of maximizing an area under the receiver operating characteristic curve is a paradigm that learns a classifier for imbalanced data (Yang and Ying, 2022). The goal is to find a classifier $\theta\in\mathbb{R}^{n}$ that maximizes the AUC score on a set of samples $\{(\mathbf{a}_{i},b_{i})\}_{i=1}^{N}$, where $\mathbf{a}_{i}\in\mathbb{R}^{n}$ and $b_{i}\in\{-1,+1\}$.
We consider the min-max formulation for AUC maximization (Ying et al., 2016; Shen et al., 2018):
@@
In Figure 2, we compare Algorithm 3 with Algorithm 2 and stochastic first-order methods. Our results evidence that Algorithm 3 outperforms SEG/SOGDA in terms of solution quality: SEG and SOGDA barely move when Algorithm 3 still makes significant progress. Compared to Algorithm 2, Algorithm 3 achieves similar solution quality eventually but require fewer samples to output an acceptable solution. It is also worth noting that Algorithm 3 exhibits (super)-linear convergence as the iterate approaches the optimal solution regardless of subsampling and inexact subproblem solving. This intriguing property was rigorously justified for the subsampled Newton method in the optimization setting (Roosta-Khorasani and Mahoney, 2019), and it would be interesting to extend such results to min-max optimization.
Refer to captionRefer to captionRefer to caption
Figure 2: Performance of all the algorithms with 3 LIBSVM datasets when $\rho=\frac{1}{N}$ is set. The numerical results are presented in terms of epoch count (Top) and computational time (Bottom).
@@ 6 Conclusion
We propose and analyze several inexact regularized Newton-type methods for finding a global saddle point of unconstrained convex-concave min-max optimization problems. Our methods are guaranteed to achieve the order-optimal iteration complexity of $O(\epsilon^{-2/3})$ and the tight complexity bound of $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$ where $\omega\approx 2.3728$ is the matrix multiplication constant. In addition, we show that our general framework and analysis on these inexact methods can lead to the first class of subsampled Newton method for solving the finite-sum min-max optimization problems with order-optimal iteration complexity. Future research directions include the extension of our methods to structured nonconvex-nonconcave min-max optimization problems and the customized implementation of our methods in real application problems.

### v5

@@ 5 Numerical Experiments
We evaluate the performance of our methods for min-max optimization problems with synthetic and real data. The baseline methods include EG and OGDA (we refer to them as EG1 and OGDA1), their stochastic variants with or without variance reduction, and second-order variants of EG and OGDA (which we refer to as EG2 and OGDA2). We exclude the methods in Huang et al. (2022) since their convergence guarantees are proved under an error bound conditions. All the methods were implemented using MATLAB R2024b on a MacBook Pro with an Intel Core i9 2.4GHz and 16GB memory.
Figure 1: Performance of all algorithms for $n\in\{50,100,200\}$ when $\rho=\frac{1}{20n}$. The numerical results are presented in terms of iteration count (Top) and computational time (Bottom).
@@ 5.1 Cubic regularized bilinear min-max problem
where $\rho>0$, the entries of $\mathbf{b}\in\mathbb{R}^{n}$ are generated independently from [1,1][-1,1] and
$A=\begin{bmatrix}1&-1&&&\\ &1&-1&&\\ &&\ddots&\ddots&\\ &&&1&-1\\ &&&&1\end{bmatrix}\in\mathbb{R}^{n\times n}.$ (5.2)
This min-max optimization problem is convex-concave and the function $f$ is $\rho$-Hessian Lipschitz. It has a unique global saddle point $\mathbf{x}^{\star}=A^{-1}\mathbf{b}$ and $\mathbf{y}^{\star}=-\frac{\rho}{2}\|\mathbf{x}^{\star}\|(A^{\top})^{-1}\mathbf{x}^{\star}$. We use the restricted gap function defined in Section 2 as the evaluation metric. In our experiment, the parameters are chosen as $\rho=\frac{1}{20n}$ and $n\in\{50,100,200\}$. Since exact second-order information is available here, we use Algorithm 2 with $\tau_{0}=0$ (the exact Hessian) and compute each subproblem with $\kappa_{m}=10^{-6}$ using our subroutine. The baseline methods include EG1, OGDA1, EG2 and OGDA2 where all of these methods require exact second-order information. We implement EG2 using the pseudocode of Bullins and Lai (2022, Algorithm 5.2 and 5.3) with the fine-tuning parameters. The implementation of OGDA2 is based on the code provided by the author of Jiang and Mokhtari (2025) with the line-search parameter $(\alpha,\beta)=(0.5,0.8)$.
Figure 1 highlights that the second-order methods can be far superior to the first-order methods in terms of solution accuracy: the first-order methods barely make any progress when the second-order methods converge successfully. Moreover, our method outperforms other second-order methods (Bullins and Lai, 2022; Jiang and Mokhtari, 2025) in terms of both iteration numbers and computational time thanks to its simple scheme without line search. Nonetheless, we remark that this does not eliminate the advantages of using line search in min-max optimization. Indeed, we find that the line-search scheme from Jiang and Mokhtari (2025) is very powerful in practice and their methods with aggressive choices of $(\alpha,\beta)$ outperforms our method in many cases. However, such choices make their method unstable. Thus, we choose $(\alpha,\beta)=(0.5,0.8)$ which is conservative yet more robust. We remark it is promising to study the line-search scheme from Jiang and Mokhtari (2025) and see if simple modifications can achieve the acceleration in a universal manner.
Figure 2: Performance of all algorithms with 3 LIBSVM datasets when $\rho=\frac{1}{N}$. The numerical results are presented in terms of iteration count (Top) and computational time (Bottom).
@@ 5.2 AUC maximization problem
The problem of maximizing an area under the receiver operating characteristic curve is a paradigm that learns a classifier for imbalanced data (Yang and Ying, 2022). The goal is to find a classifier $\theta\in\mathbb{R}^{n}$ that maximizes the AUC score on a set of samples $\{(\mathbf{a}_{i},b_{i})\}_{i=1}^{N}$, where $\mathbf{a}_{i}\in\mathbb{R}^{n}$ and $b_{i}\in\{-1,+1\}$. The min-max formulation for AUC maximization (Ying et al., 2016; Shen et al., 2018) is:
@@
Figure 2 shows that Algorithm 3 outperforms stochastic first-order methods in terms of solution quality: stochastic first-order methods barely move when Algorithm 3 makes significant progress. Notably, Algorithm 3 exhibits (super)-linear convergence as the iterates approach the optimal set regardless of subsampling and inexact subproblem solving. This intriguing property has been rigorously justified for subsampled Newton methods for convex optimization (Roosta-Khorasani and Mahoney, 2019), and extending such results to convex-concave min-max optimization is an interesting future direction.
@@ 6 Conclusion
We propose several inexact regularized Newton-type methods for finding a global saddle point of unconstrained convex-concave min-max optimization problems. Our inexact methods are guaranteed to achieve an iteration complexity of $O(\epsilon^{-2/3})$ and an improved complexity bound of $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$ where $\omega\approx 2.3728$ is the matrix multiplication constant. In addition, we show that our general framework and analysis yields the first class of subsampled Newton method for solving finite-sum min-max optimization problems with the same iteration complexity of $O(\epsilon^{-2/3})$. Future research directions include the extension of our methods to structured nonconvex-nonconcave min-max optimization problems and the customized implementation of our methods in real application problems.